40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 913     Entry time: Fri Apr 13 13:35:37 2012
Author: Tara, Frank 
Type: Summary 
Category: NoiseBudget 
Subject: PDH readout noise level verification 

we re-checked the electronic noise level for each PDH loop and also checked how the noise level rises with a shot noise limited light source (halogen bulb) using the exact same setup as we use daily.
Before we had only checked the individual parts, e.g. the shot noise limit of our RF-PD, not including the rest of our setup (mixer, LO or filter stages). This was just to confirm that nothing is broken and our PDs are still OK.

Measurements were taken right after the mixer / LP-filter, but still connected to PDH servo. This is OUT1 of the common path of the TTFSS-box  and the input on the front of our ACAV PDH servo. We blocked the laser and adjusted the power of the bulb and measured the electronic noise floor using the SR785. The dark noise spectrum is the one already shown in the last noise budget. Below the plots of noise level, measured around 500Hz (100 avgs, wide marker) vs different photocurrents of the PDs. DC impedance is 2kOhms, regular DC level for RCAV is 1.7V and 1.47V for ACAV. RF-impedance is the same as measured some time ago for those diodes. Data in folder 4/11/2012 on svn.

SN_RCAV.png  SN_ACAV.png

Conclusion: New measurements agree very well with the old measurements. So our PDs still work fine and there is no excess noise in our readout we missed.

ELOG V3.1.3-